9s-5g=-4u,for s. This is solving equations and formulas. Trying to help son and don't understand.
Answer by nerdybill(7384)

9s-5g=-4u,for s
The idea is to isolated the interested variable (in this case, it's 's') to one side of the equation.
.
Starting with:
9s-5g=-4u
Add 5g to both sides:
9s-5g+5g = -4u+5g
9s = -4u+5g
Now, divide both sides by 9:
(9s)/9 = (-4u+5g)/9
s = (-4/9)u + (5/9)g (This is what they're looking for)